OpenStudy (anonymous):

Partial fraction decomposition.... Again, I went wrong somewhere. Not sure why I am not getting this... Partial fraction decomposition: (5x^2+20x+6)/x^3+2x^2+x Here's what I got but I think it is wrong: (5x^2+20x+6)/x(x^2+2x+1) (5x^2+20x+6)/x(x+1)^2 (A/x)+(B/(x+1))+(C/(x+1)^2) A(x+1)(x+1)^2 + B(x)(x+1)^2 + C(x)(x+1) (5x^2+20x+6)=Ax^3+Bx^3+3Ax^2+2Bx^2+Cx^2+3Ax+Bx+Cx+A (5x^2+20x+6)=x^3(A+B) + x^2(3A+2B+C) + x(3A+B+C) + A A+B=0 3A+2B+C=5 3A+B+C=20 A=6 A=6 B=-6 C=-1 But this isn't right. Where did I go wrong?

OpenStudy (anonymous):

why do you expand them for starters?

OpenStudy (anonymous):

(5x^2+20x+6)/x(x+1)^2=(A/x)+(B/(x+1))+(C/(x+1)^2) (5x^2 +20x +6 = A(x+1)^2 +Bx(x+1) + C(x)

OpenStudy (anonymous):

u made alot of mistakes :|

OpenStudy (anonymous):

^% thats correct

OpenStudy (anonymous):

sub x=0: 6 = A A+6 .

OpenStudy (anonymous):

sub x=-1 : 5 -20+6 = -C C= 9

OpenStudy (anonymous):

sub x=1 : 31= 4A+2B+C B = (31-4A-C) / 2 = (31-24 -9) / 2 = -1

OpenStudy (anonymous):

A=6, B=-1, C=9

OpenStudy (anonymous):

comparing coeffiecents is a very long method and bound to make tons of mistakes